OpenStudy (anonymous):

The least common multiple of 54c2de3 and 38c4 is _____. The least common multiple of 54c2de3 and 38c4 is _____. @Mathematics

OpenStudy (anonymous):

the numbers are exponets

OpenStudy (anonymous):

exept for 54 and 38

OpenStudy (turingtest):

\[54c^2de^3\] and \[38c^4\] the real numbers coefficients are\[ 54=2*3^3\]and\[38=2*19\]so they only have a two in common \[2*3^2*19*c^4*d*e^3=342c^4de^3\]

OpenStudy (turingtest):

last line should be\[2∗3^3∗19∗c^4∗d∗e^3=1026c^4de^3\]sorry...